Given f(x)= 2x-1 x<3
-2 x > or = 3
Evaluate f(1)
Evaluate f(3)
Evaluate f(5)

Answer by stanbon(48535)

Given f(x)= 2x-1 x<3
-2 x > or = 3
Evaluate f(1)
Since 1<3, f(1)=2(1)-1=1
--------------
Evaluate f(3)
Since 3=3, f(3)=-2
----------
Evaluate f(5)
Since 5>3, f(5)=2(5)-1=9
-------------
Cheers,
Stan H.